OpenStudy (anonymous):

Can someone tell me if these are right 21. Solve 1/3x+9 - 2/x+3 = 2 x=1 x=-23/6 x=-5/2******* x=-5 Solve 8/x + 15/x^2= -1 x=5,-3***** x=-5,-3 x=-5,3 x=5,3

jimthompson5910 (jim_thompson5910):

both are incorrect unfortunately

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

can you show me how you knew that i dont want the answer i just need help of ehy i got em right

jimthompson5910 (jim_thompson5910):

It might help to multiply everything by the LCD (x+3)(3x+9) to clear out the fractions 1/(3x+9) - 2/(x+3) = 2 x+3 - 2(3x+9) = 2(x+3)(3x+9) now solve that second equation for x

OpenStudy (anonymous):

x+3-2(3x^3)= 2(x+3)(3x^3)

jimthompson5910 (jim_thompson5910):

no, 1/(3x+9) - 2/(x+3) = 2 x+3 - 2(3x+9) = 2(x+3)(3x+9) x+3 - 6x - 18 = 2(3x^2 + 18x + 27) x+3 - 6x - 18 = 6x^2 + 36x + 54 -5x - 15 = 6x^2 + 36x + 54 0 = 6x^2 + 36x + 54 + 5x + 15 0 = 6x^2 + 41x + 69 6x^2 + 41x + 69 = 0 I'll let you finish

OpenStudy (anonymous):

23/6

jimthompson5910 (jim_thompson5910):

close, very close

OpenStudy (anonymous):

That the answer

jimthompson5910 (jim_thompson5910):

close to it, yes

OpenStudy (anonymous):

im correct:)

jimthompson5910 (jim_thompson5910):

i think you meant to say -23/6 instead of 23/6

jimthompson5910 (jim_thompson5910):

so yes, you are correct

OpenStudy (anonymous):

& the second one is B

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you so much

jimthompson5910 (jim_thompson5910):

np